given: f(x)=(1/4)^x. how do I write down the equation of f^-1(x) in the form y=...
Answer by Theo(13342)

start with f(x) = (1/4)^x.

set y = f(x) and the equation becomes y = (1/4)^x.

replace y with x and x with y and the equation becomes x = (1/4)^y.

now you want to solve for y.

x = (1/4)^y if and only if log(1/4)(x) = y

your equation becomes y = log(1/4)(x)

set y = f-1(x) and the equation becomes f-1(x) = log(1/4)(x)

the graph of both equations is shown below.
since they are inverse equations, they are reflections about the equation of y = x.

$$$

they are inverse equations if (x,y) in f(x) is equal to (y,x) in f-1(x).

the intersection of the both graphs with a line perpendicular to the line y = x shows this to be true.

the red curve is the normal equation and the blue curve is the inverse equation.

the reflection of the point (-1.448,7.448) in the normal equation is the point (7.448,-1.448) in the inverse equation, confirming that they are inverse equations.


in general, if you start with y = bx, and you want to find the inverse equation, then replace x with y and y with x to get the equation x = by

this is true if and only if y = logb(x)

the inverse equation to y = bx is therefore y = logb(x).

here's a nice reference on logarithmix and exponential functions you might find useful.

http://www.themathpage.com/aprecalc/logarithmic-exponential-functions.htm